# Jaipur-302017, INDIA

### v2

@@ Abstract
We investigate the large $N$ limit of a complex action unitary matrix model with Fisher-Hartwig singularities, motivated by QCD-inspired models with complexified potentials. We show that the model exhibits multiple ungapped phases and a single gapped phase. The phases are characterized by Fisher-Hartwig singularities in the complex plane. We show that the phase transitions are third order, with transitions between ungapped phases forbidden. We also briefly discuss the implications for the QCD phase diagram at the end.
1 Introduction
@@
The study of spectral and critical properties of random matrix models has become a central interdisciplinary topic, as these models naturally arise in quantum field theory, condensed matter physics, number theory, and statistics 27; 23; 5; T. Guhr, A. Müller–Groeling, and H. A. Weidenmüller (1998). Unitary matrix models appear in a wide variety of physical systems, such as gauge theories Gross and Witten (1980); Aharony et al. (2004); Dumitru et al. (2005); Hands et al. (2010), two-dimensional Quantum Gravity Brézin and Kazakov (1990); Douglas and Shenker (1990); Gross and Migdal (1990), the Ising model and disordered mesoscopic conductors McCoy and Wu (1973); Muttalib et al. (1993); Deift et al. (2013), as well as topological field and string theories Vleeshouwers and Gritsev (2021). They are also connected to the Painlevé equations and their $\tau$-functions J. Forrester and S. Tierz (2001); Szabo and Tierz (2012). Simpler/limiting versions of unitary matrix models have been investigated for decades by both mathematicians and physicists; see Refs. Mandal (1990); Russo (2020); Santilli and Tierz (2020, 2022) and the references therein. We also refer to the excellent literature reviews Refs. Rossi et al. (1998) and Marino (2005).
@@
Depending on the system, the matrix model description may be non-perturbative, which has its own obvious advantages. In fact, in certain cases the non-perturbative structure can be determined in complete detail Ahmed and Dunne (2017, 2018), and in such cases matrix models serve as a wonderful test lab to explore resurgent asymptotics. Additionally, matrix models have attracted attention in recent years due to their relevance to systems that exhibit the sign problem, which obstructs the use of standard techniques such as Monte-Carlo in lattice gauge theory Hands (2007). Recent progress in resolving the sign problem has involved methods such as complex Langevin dynamics Aarts et al. (2010); Boguslavski et al. (2025); Klauder (1984) and the Lefschetz-thimble approach Fujii et al. (2015); Cristoforetti et al. (2013); Nishimura et al. (2024).
@@
SU(NN) matrix models are a subset of unitary matrix models, where the integral which defines the partition function is restricted to SU($N$) matrices. We write
$Z=\int_{\text{SU}(N)}dU\,e^{-\mathcal{S}(U)}$ (1)
@@
where $\mathcal{S}$ is the action. Expectation values in the model are defined by
$\langle f(U)\rangle=\int_{\text{SU}(N)}dU\,f(U)\,e^{-\mathcal{S}(U)}.$ (2)
In this work we focus on the action $\mathcal{S}=N\,\text{Tr}\,V(U)$ with the potential
@@
$V(z)=-\sigma\left(\ln(1+\alpha z)+\ln\left(1+\frac{1}{\gamma z}\right)\right).$ (3)
This potential is complex valued. We note here that a large number of unitary models studied in the literature have a real potential. As the potential (3) is complex and the model (1) is an SU($N$) model, the relation $\langle U\rangle=\langle U^{-1}\rangle$ may not necessarily hold for our model, whereas in the case of U($N$) models with real potentials it always holds. The branch points of the logarithm in (3) are Fisher-Hartwig singularities (see Ref. Fahs (2021) and the references therein), and their existence implies a rich phase structure.
The $\gamma\to\infty$ limit of this model has been studied in Ref. Hands et al. (2010), where it describes the statistics of a two-level QCD (Quantum Chromodynamics) system restricted to a small sphere. In the triple-scaling limit $\sigma\to 0,\alpha\to 0,\gamma\to\infty$ while keeping $\alpha\sigma$ and $\sigma/\gamma$ equal and constant, the model reduces to the SU($N$) version of the well-known Gross-Witten-Wadia model Gross and Witten (1980). Thus our model can be viewed as a generalization of several existing models.
This work is organized as follows. In Sec. 2, we summarize the standard approach used to study unitary matrix models, and introduce our notation. In Sec. 3, we discuss the ungapped phase and identify four distinct ungapped phases, each characterized by the location of poles of the potential. In Sec. 4, we examine the gapped phase. In Sec. 5 we briefly discuss the transitions between these phases. Finally, in Sec. 6, we summarize our main results and outline possible extensions of this work. We restrict our work to the large $N$ limit, as there are no phase transitions at finite $N$. We also restrict the coupling $\sigma$ and the parameters $\alpha,\gamma$ to positive values.
2 Formalism
The partition function (1) is an integral over SU($N$) matrices; that is, the integral is constrained to those unitary matrices $U$ which satisfy $\det U=1$. This constraint can also be written as $\text{Tr}\,\ln U=0$. We can convert the constrained SU($N$) integral to an unconstrained U($N$) integral by introducing a Lagrange multiplier $\mathscr{N}$. In the large $N$ limit, this amounts to modifying the potential as
$V(z)\to V(z)+\mathscr{N}\ln z,\qquad\mathscr{N}\in\mathbb{R}.$ (4)
@@ 2 Formalism
$Z=\int\Big(\prod_{i=1}^{N}d\theta_{i}\Big)\,e^{-\mathcal{S}_{\text{eff}}},$ (5)
@@
$\mathcal{S}_{\text{eff}}=N\sum_{i=1}^{N}V(e^{i\theta_{i}})-\frac{1}{2}\sum_{i,j=1}^{N}\ln{\sin^{2}{\left(\frac{\theta_{i}-\theta_{j}}{2}\right)}}+iN\mathscr{N}\sum_{i=1}^{N}\theta_{i}.$ (6)
The second sum in the above expression is the Vandermonde contribution. In this form, an approximation to the integral can be found by using the saddle point method. However, since we are working in the exact large $N$ limit, this approximation is actually exact. The definition of saddle points, $\partial S_{\text{eff}}/\partial\theta_{i}=0$ for every $i$, gives
$V^{\prime}(\theta_{i})+i\mathscr{N}=\frac{1}{N}\sum_{i\neq j}\cot{\left(\frac{\theta_{i}-\theta_{j}}{2}\right)}.$ (7)
It is clear from the left hand side of the equation that the saddles may be complex.
@@
$\bar{\rho}(\theta)=\frac{1}{N}\sum_{i}\delta(\theta-\theta_{i}),$ (8)
which allows the following replacement when taking the large $N$ limit:
$\frac{1}{N}\sum_{i}^{N}f(\theta_{i})\to\int d\theta\,\bar{\rho}(\theta)f(\theta).$ (9)
@@
$z_{i}V^{\prime}(z_{i})+\mathscr{N}=\frac{1}{N}\sum_{i\neq j}\frac{z_{i}+z_{j}}{z_{i}-z_{j}}$ (10)
and its large $N$ limit is
$zV^{\prime}(z)+\mathscr{N}=\mathscr{P}\int_{\mathcal{C}}\frac{dz^{\prime}}{2\pi iz^{\prime}}\,\rho(z^{\prime})\,\frac{z+z^{\prime}}{z-z^{\prime}},\qquad z\in\mathcal{C}.$ (11)
In the above relation $\mathscr{P}$ denotes the principal value of the integral and $\rho(z):=2\pi\bar{\rho}(\theta)$ for all $z\in\mathcal{C}$, where $\mathcal{C}$ is the contour in the complex-eigenvalue space on which the saddle points coalesce (also called the support of $\rho(z)$). The density of the saddle points on $\mathcal{C}$ is
$D(z)=\dfrac{\rho(z)}{2\pi iz}.$ (12)
@@
The problem now is to find the spectral function $\rho(z)$ and its support $\mathcal{C}$ given the potential (3). This is a Reimann-Hilbert problem Brézin et al. (1978); Marino (2005), and it can be solved in a rather straight-forward manner if we restrict to solutions in which $\mathcal{C}$ is a single continuous curve. Such a solution is called a single cut solution. If the potential is real and model is a U($N$) matrix model (i.e. $\mathcal{N}=0$), then the saddle points are found exclusively on the unit circle and the support $\mathcal{C}$ is simply the unit circle or an arc of a unit circle. In the present case the saddle points may be found off the unit circle.
The support $\mathcal{C}$ can be found by solving the differential equation
$\dfrac{dz}{d\theta}=\frac{iz}{\rho(z)},\qquad z\equiv z(\theta).$ (14)
The expression for the expectation value (2) simplifies in the large $N$ limit to
$\langle f(U)\rangle=\int_{C}\dfrac{dz}{2\pi iz}\rho(z)f(z).$ (15)
@@
$\det U=1\implies\sum_{i}\theta_{i}=0\implies\int_{C}\frac{dz}{2\pi iz}\rho(z)\ln{z}=0.$ (16)
In the next section, we discuss the ungapped phases of the model.
3 Ungapped phases
Figure 1: Left: support of the spectral function in the ungapped phase A (small $\alpha$ large $\gamma$ phase) for a few values of $\alpha$. The dot at $z=-1/\gamma$ is the pole which is inside $\mathcal{C}$, while the other pole (not shown) is outside. Right: phase boundary between the ungapped phase A and the gapped phase.
Figure 2: Support of the spectral function (left panels) and the phase boundary between the gapped and ungapped phases (right panels). The panels from top to bottom correspond to ungapped phases B, C and D in the same order.
@@ 3 Ungapped phases
$\sigma\left(-1+\dfrac{1}{1+\alpha z}+\dfrac{1}{1+\gamma z}\right)+\mathscr{N}=\mathscr{P}\oint_{\mathcal{C}}\frac{dz^{\prime}}{2\pi iz^{\prime}}\,\rho(z^{\prime})\,\frac{z+z^{\prime}}{z-z^{\prime}},\qquad z\in\mathcal{C},$ (17)
become closed contour integrals. Thus, we can immediately utilize the standard theorems of elementary Complex Analysis to solve the model and compute various quantities of interest.
We assume the following ansatz based on the poles of the potential ($z=-\frac{1}{\alpha},-\frac{1}{\gamma}$):
$\rho(z)=\sum_{n=0}^{\infty}\frac{A_{n}}{z^{n}}+\sum_{n=0}^{\infty}\frac{B_{n}}{(z+1/\alpha)^{n}}+\sum_{n=0}^{\infty}\frac{C_{n}}{(z+1/\gamma)^{n}}.$ (18)
It is necessary to include a pole at $z=0$ in the ansatz in order to satisfy the normalization condition (13), and this pole must always be inside $\mathcal{C}$. The coefficients in the ansatz are determined by substituting it in (17) and comparing both sides. The normalization condition can be used to fix any one coefficient as well.
There are four different solutions for $\rho(z)$ depending on whether the poles $z=-\frac{1}{\alpha}$ and $z=-\frac{1}{\gamma}$ are inside or outside $\mathcal{C}$. Each of these solutions corresponds to a distinct ungapped phase. We use the following nomenclature:
@@
$\rho(z)=1+\sigma+\frac{\sigma}{1+\gamma z}-\frac{\sigma}{1+\alpha z}.$ (19)
The comparison also shows that the Lagrange multiplier $\mathcal{N}$ vanishes. With the spectral function known, the SU($N$) constraint (16) simplifies to
$(1+\sigma)\ln{r_{0}}+\sigma\ln{\left(\frac{1-r_{0}\alpha}{r_{0}\gamma-1}\right)}+\sigma\ln\gamma=0$ (20)
@@
$\alpha|z|^{1+1/\sigma}=\left|z+\frac{1}{\alpha}\right|^{-1}\,\left|z+\frac{1}{\gamma}\right|.$ (21)
Just as expected, $\mathcal{C}$ is not a unit circle in this phase. See the left panel in Fig. 1 for plots of $\mathcal{C}$ in phase $A$.
@@
$\alpha=\frac{\gamma r_{0}-\sigma-1}{r_{0}(\gamma r_{0}(\sigma+1)-2\sigma-1)}$ (22)
@@
$W_{n}:=\frac{1}{N}\langle\mathrm{tr}\,U^{n}\rangle=\oint_{C}\dfrac{dz}{2\pi iz}\rho(z)z^{n}=\begin{cases}(-1)^{n+1}\frac{\sigma}{\gamma^{n}}\hskip 42.67912ptn\geq 1\\ (-1)^{n+1}\sigma\alpha^{|n|}\hskip 34.14322ptn\leq-1\end{cases}$ (23)
If our model results from a lattice gauge theory, then these moments are related to the winding Wilson loops over a single plaquette Okuyama (2017). We observe that
$W_{n}\neq W_{-n}$ (24)
which indicates the presence of the sign problem. This behavior is completely expected, as our model’s potential is complex, in addition to having an SU($N$) geometry. In the Gross-Witten-Wadia model Gross and Witten (1980), the single-winding Wilson loop ($W_{1}$) is related the string tension $\tau$ and the lattice spacing $a$ as $W_{1}=e^{-a^{2}\tau}$. If we assume that this continues to hold if our model is connected to a lattice gauge theory, then the beta function of $1/\sigma$ has the expression
$\beta=\dfrac{2W_{1}}{\sigma^{2}}\dfrac{\ln W_{1}}{\partial_{\sigma}W_{1}}.$ (25)
For phase A this yields $\beta=\dfrac{2}{\sigma}\ln{\dfrac{\sigma}{\gamma}}$.
Lastly, we look at the free energy, defined as
$F=-\frac{1}{N^{2}}\ln Z.$ (26)
As the free energy in the large $N$ limit is just the rescaled effective action, $F=\frac{1}{N^{2}}S_{\text{eff}}$, the continuum version of (6) can be used to directly calculate the free energy. Another method is to first find the derivatives of the free energy with respect to each parameter of the model using
$\frac{\partial F}{\partial t}=\left\langle\frac{\partial V(U)}{\partial t}\right\rangle=\oint_{C}\frac{dz}{2\pi iz}\rho(z)\frac{\partial V(z)}{\partial t},\qquad t=\sigma,\alpha,\gamma,$ (27)
@@
$\frac{\partial F}{\partial\alpha}=\frac{\sigma^{2}}{\alpha-\gamma},\qquad\frac{\partial F}{\partial\gamma}=-\frac{\sigma^{2}}{\gamma}+\frac{\sigma^{2}}{\gamma-\alpha},\qquad\frac{\partial F}{\partial\sigma}=2\sigma\ln{\left(1-\frac{\alpha}{\gamma}\right)}.$ (28)
@@
$F=\sigma^{2}\ln\left(1-\frac{\alpha}{\gamma}\right).$ (29)
In the tables below we summarize the important quantities in each phase. See Figs. 1 and 2 for plots of the support of the spectral function and the phase diagram near the ungapped-gapped transition for each ungapped phase.
@@
Beta function $\dfrac{2}{\sigma}\ln{\dfrac{\sigma}{\gamma}}$ $\dfrac{2}{\sigma}\ln{\left(\dfrac{\sigma}{\alpha}+\dfrac{\sigma}{\gamma}\right)}$
@@
Beta function $\dfrac{2}{\sigma}\ln{\dfrac{\sigma}{\alpha}}$ Undefined

### v3

# Spectral and Phase Structure of a Unitary Matrix Model with Fisher-Hartwig Singularities

@@ Abstract.
We investigate a unitary matrix model with a complex potential with Fisher-Hartwig singularities. We show that the model exhibits finite-$N$ phase transitions. The order of the phase transition is coupling-dependent. At large-$N$, these transitions are replaced by third order Gross-Witten-Wadia transitions between multiple ungapped phases and a single gapped phase, with transitions between ungapped phases forbidden. At both finite and large $N$, the phases are characterized by the locations of the Fisher-Hartwig singularities in the complex plane.
1. Introduction
@@
The study of spectral and critical properties of random matrix models has become a central interdisciplinary topic, as these models naturally arise in quantum field theory, condensed matter physics, number theory, and statistics [28, 14, 5, 19]. Unitary matrix models appear in a wide variety of physical systems, such as gauge theories [18, 2, 13, 21], two-dimensional Quantum Gravity [9, 12, 17], the Ising model and disordered mesoscopic conductors [27, 30, 11], as well as topological field and string theories [38]. They are also connected to the Painlevé equations and their $\tau$-functions [22, 37]. Simpler/limiting versions of unitary matrix models have been investigated for decades by both mathematicians and physicists; see Refs. [24, 34, 35, 36] and the references therein. We also refer to the excellent literature reviews Refs. [33] and [25].
@@
Depending on the system, the matrix model description may be non-perturbative, which has its own obvious advantages. Matrix models can serve as wonderful test labs to explore resurgent asymptotics. In fact, in certain cases the non-perturbative structure can be determined in complete detail [26, 3, 4]. Additionally, matrix models have attracted attention in recent years due to their relevance to systems that exhibit the sign problem, which obstructs the use of standard techniques such as Monte-Carlo in lattice gauge theory [20]. Recent progress in resolving the sign problem has involved methods such as complex Langevin dynamics [1, 6, 23] and the Lefschetz-thimble approach [15, 10, 31].
@@
SU(NN) matrix models are a subset of unitary matrix models, where the integral which defines the partition function is restricted to SU($N$) matrices. In standard notation, the partition function is
(1) $Z=\int_{\text{SU}(N)}dU\,e^{-\mathcal{S}(U)/g}$
@@
where $\mathcal{S}$ is the action and $g$ is the coupling. We will also use the variable $\tau=1/g$ when convenient. Expectation values in the model are defined by
(2) $\langle h(U)\rangle=\frac{1}{Z}\int_{\text{SU}(N)}dU\,h(U)\,e^{-\mathcal{S}(U)/g}.$
In this work we focus on the action
@@
(3) $\mathcal{S}(U)=-\text{Tr}\,\left(\ln(1+\alpha U)+\ln\left(1+\frac{1}{\gamma}U^{\dagger}\right)\right),\qquad\alpha,\gamma>0.$
As this SU($N$) model has a complex action, we expect it to violate the relation $\langle U\rangle=\langle U^{\dagger}\rangle$, which always holds for U($N$) models with real actions. A more interesting feature of the action is the presence of Fisher-Hartwig singularities, and their existence implies a rich phase structure.
A generalized version of this model describes the partition function of a weakly coupled gauge theory [2], and its low temperature limit (which corresponds to $\gamma\to\infty$) has been studied in Ref. [21]. A U($N$) matrix model with an action similar to (3) famously describes the Ising spin correlation function [39]. In the triple-scaling limit $g\to 0,\alpha\to 0,\gamma\to\infty$ while keeping $g/\alpha$ and $g\gamma$ equal and constant, the action (3)‘ reduces to the well-known Gross-Witten-Wadia action [18].
2. Behaviour at finite NN
The partition function (1) is defined as an integral over SU($N$) matrices; that is, the integral is restricted to unitary matrices $U$ satisfying $\det U=1\iff\text{Tr}\,\ln U=0$. The SU($N$) integral can be to an unconstrained U($N$) integral by means of a Fourier decomposition.
(4) $\int_{\text{SU}(N)}dU\,e^{-\mathcal{S}(U)/g}=\frac{1}{2\pi}\sum_{k=-\infty}^{\infty}\int_{\text{U}(N)}dU\,e^{-\mathcal{S}(U)/g+i\,k\,\text{Tr}\,\ln U}.$
The SU($N$) problem is now a U($N$) problem, but at the cost of an additional infinite sum at the end.
At finite-$N$, the U($N$) integrals may be computed by using the standard Toeplitz determinant representation
(5) $\int_{\text{U}(N)}\,\det f(U)=\det(f_{i-j})_{i,j=1,\dots,N}$
where $f_{n}$ are the Fourier coefficients of the symbol $f(z)$
(6) $f_{n}=\oint_{|z|=1}\dfrac{dz}{2\pi i}\dfrac{f(z)}{z^{n+1}}.$
Combining this with (4), gives the following representation for the SU($N$) integral:
(7) $\int_{\text{SU}(N)}dU\,e^{-\mathcal{S}(U)/g}=\frac{1}{2\pi}\sum_{k=-\infty}^{\infty}\det(f_{i-j+k})_{i,j=1,\dots,N}$
With the partition function known, the free energy can be calculated as
(8) $F=-\dfrac{1}{N^{2}}\ln Z$
The symbol associated with the action (3) is
(9) $f(z)=(1+\alpha z)^{\tau}\left(1+\dfrac{1}{\gamma z}\right)^{\tau},\qquad\tau=1/g.$
When $\tau$ is not an integer, the points $z=-1/\alpha$ and $z=-1/\gamma$ become branch points. If these branch points lie on the unit circle, then these are called the Fisher-Hartwig singularities. Thus there are four parameter regimes, defined by the positions of the branch points relative to the unit circle. As the parameter values change and the branch points cross the unit circle, the Fourier coefficients pick up contributions from the branch cuts, causing the partition function, and consequently the free energy, to have discontinuous derivatives. Of course, if $\tau\in\mathbb{Z}$ then the partition function is a smooth function everywhere. We restrict the discussion to $\gamma>1$ and $0<\alpha<\gamma$, as these parameter ranges are sufficient to study the behaviour described above.
Consider first the small $\alpha$ case ($0<\alpha<1<\gamma$). The branch point $z=-1/\alpha$ lies outside the unit circle while $z=-1/\gamma$ lies inside it, so no branch cut line overlaps the unit circle. The Fourier coefficients can be evaluated explicity in terms of the Gauss hypergeometric function using standard methods:
(10) $f^{\text{small}}_{n}=u_{n}(\alpha,\gamma)\begin{pmatrix}\tau\\ |n|\end{pmatrix}\,_{2}F_{1}(|n|-\tau,-\tau;1+|n|;\alpha/\gamma)$
where
(11) $u_{n}(\alpha,\gamma)=\begin{cases}\alpha^{n},\qquad n\geq 0\\ \gamma^{n},\qquad n\leq 0.\end{cases}$
As the Gauss hypergeometric function is analytic, $f^{\text{small}}_{n}$ are smooth functions of $\alpha$ for all $\alpha\in[0,1]$. See Ref. [7] for an alternative expression for the small-$\alpha$ Fourier coefficients and a representation of the partition function as a Fredholm determinant.
When $\alpha$ becomes large ($1<\alpha<\gamma$), the branch point $z=-1/\alpha$ lies inside the unit circle, and the branch cut of $(1+\alpha z)^{\tau}$ intersects the unit circle. This time the Fourier coefficients pick up an additional contribution from the branch cut. The coefficients are significantly more complicated and are given by
(12) $f^{\text{large}}_{n}=f^{\text{small}}_{n}+\dfrac{(-1)^{n}}{\pi}\dfrac{\sin(\pi\tau)}{\tau+1}\alpha^{n}(\alpha-1)^{\tau+1}\left(1-\dfrac{\alpha}{\gamma}\right)^{\tau}\,A(n,\tau,\alpha,\gamma)$
where $A(n,\tau,\alpha,\gamma)$ is shorthand for the Appell function
(13) $A(n,\tau,\alpha,\gamma):=F_{1}\left(\tau+1;n+\tau+1,-\tau;\tau+2;1-\alpha,\dfrac{1-\alpha}{1-\alpha/\gamma}\right).$
The term containing this function is the contribution from the branch cut, and it vanishes if $\tau\in\mathbb{Z}$. While the Appell function is infinitely differentiable near $\alpha=1$, the pre-factor is not. The behaviour of the branch cut contribution near $\alpha=1$ is
(14) $(-1)^{n}\dfrac{\sin(\pi\tau)}{\pi(1+\tau)}\,\left(1-1/\gamma\right)^{\tau}\,\epsilon^{\tau+1},\qquad\epsilon=\alpha-1.$
It is clear that there is a discontinuity in the $k$-th derivative if $\tau\neq\mathbb{Z}$, where $k=1+\mathrm{ceiling}(\tau)$, and this discontinuity propagates to the partition function and free energy. Therefore, in the Ehrenfest classification, the order of the phase transition at $\alpha=1$ is $1+\mathrm{ceiling}(\tau)$.
With the Fourier coefficients calculated, the finite-$N$ partition function and free energy can be studied. We leave the full expressions of these quantities for the SU($N$) model for a future work. Fig. 1 shows the second derivative of the $N=3$ free energy of the U($N$) model with respect to $\alpha$. It is clear that for $\tau=0.5$ and $1.5$, the order of transition is $2$ and $3$ respectively, consistent with the discussion above.
In the large-$N$ limit, the coupling $g$ must be scaled inversely with $N$, otherwise the model reduces to the trivial model $Z=\int dU$. Thus, $\tau\propto N$ in the large-$N$ limit, and consequently, the order of the phase transition approaches infinity. This indicates that the large-$N$ partition function becomes smooth at $\alpha=1$ (and also $\gamma=1$). In other words, the finite-$N$ phase transition disappears in the large-$N$ limit. In the next section we work in the exact large-$N$ limit and show that this is indeed the case.
Figure 1. Second derivative of the free energy of the U($N$) model with respect to $\alpha$ at $N=3$.
3. Large-$N$ limit
We work in the double scaling limit $N\to\infty,g\to 0$ while keeping constant the following combination
(15) $\sigma=\dfrac{1}{Ng}.$
We define the potential $V$ by
(16) $\mathcal{S}(U)/g=N\,\mathrm{Tr}\,V(U).$
The potential corresponding to the action (3) is
(17) $V(z)=-\sigma\left(\ln(1+\alpha z)+\ln\left(1+\frac{1}{\gamma z}\right)\right).$
In the exact large-$N$ limit the SU($N$) integral with potential $V(z)$ is equivalent to a U($N$) integral with the potential $V(z)+\mathscr{N}\ln z$ [21]. Here $\mathscr{N}$ is a Lagrange multiplier that enforces the SU($N$) constraint $\det U=1$, and it must be determined self-consistently by demanding $\langle\mathrm{Tr}\,\ln U\rangle=0$.
@@ 3. Large-$N$ limit
(18) $Z=\int\Big(\prod_{i=1}^{N}d\theta_{i}\Big)\,e^{-\mathcal{S}_{\text{eff}}/g},$
@@
(19) $\mathcal{S}_{\text{eff}}/g=N\sum_{i=1}^{N}V(e^{i\theta_{i}})-\frac{1}{2}\sum_{i,j=1}^{N}\ln{\sin^{2}{\left(\frac{\theta_{i}-\theta_{j}}{2}\right)}}+iN\mathscr{N}\sum_{i=1}^{N}\theta_{i}.$
The second sum in the above expression is the Vandermonde contribution. In this form, an approximation to the integral can be found by using the saddle point method. As we are working in the exact large-$N$ limit, this approximation is actually exact. The saddle point equation, $\partial\mathcal{S}_{\text{eff}}/\partial\theta_{i}=0$ for every $i$, works out to be
(20) $V^{\prime}(\theta_{i})+i\mathscr{N}=\frac{1}{N}\sum_{i\neq j}\cot{\left(\frac{\theta_{i}-\theta_{j}}{2}\right)}.$
It is clear from the form of the equation that the saddles may be complex.
@@
(21) $\bar{\rho}(\theta)=\frac{1}{N}\sum_{i}\delta(\theta-\theta_{i}),$
which allows the following replacement when taking the large-$N$ limit:
(22) $\frac{1}{N}\sum_{i}^{N}f(\theta_{i})\to\int d\theta\,\bar{\rho}(\theta)f(\theta).$
@@
(23) $z_{i}V^{\prime}(z_{i})+\mathscr{N}=\frac{1}{N}\sum_{i\neq j}\frac{z_{i}+z_{j}}{z_{i}-z_{j}}$
and its large-$N$ limit is
(24) $zV^{\prime}(z)+\mathscr{N}=\mathscr{P}\int_{\mathcal{C}}\frac{dz^{\prime}}{2\pi iz^{\prime}}\,\rho(z^{\prime})\,\frac{z+z^{\prime}}{z-z^{\prime}},\qquad z\in\mathcal{C}.$
In the above relation $\mathscr{P}$ denotes the principal value of the integral and $\rho(z):=2\pi\bar{\rho}(\theta)$ for all $z\in\mathcal{C}$, where $\mathcal{C}$ is the contour on which the saddle points coalesce (also called the support of $\rho(z)$). The density of the saddle points on $\mathcal{C}$ is
(25) $D(z)=\dfrac{\rho(z)}{2\pi iz},\qquad z\in\mathcal{C}.$
@@
The problem now is to find the spectral function $\rho(z)$ and its support $\mathcal{C}$ given the potential (17). This is a Reimann-Hilbert problem [8, 25], and it can be solved in a rather straight-forward manner if we restrict to solutions in which $\mathcal{C}$ is a single continuous curve. Such a solution is called a single cut solution. For U($N$) matrix integrals with real action the saddle points are found exclusively on the unit circle and $\mathcal{C}$ is simply the unit circle or an arc of a unit circle. As the present model is an SU($N$) integral with a complex action, the saddle points may be found off the unit circle. Once $\rho(z)$ is known, $\mathcal{C}$ can be found by solving
(27) $\dfrac{dz}{d\theta}=\frac{iz}{\rho(z)},\qquad z\equiv z(\theta).$
The expression for the expectation values simplifies drastically in the large-$N$ limit:
(28) $\frac{1}{N}\,\langle\mathrm{Tr}\,h(U)\rangle\to\int_{C}\dfrac{dz}{2\pi iz}\rho(z)\,h(z).$
@@
(29) $\det U=1\implies\sum_{i}\theta_{i}=0\implies\int_{C}\frac{dz}{2\pi iz}\rho(z)\ln{z}=0.$
Next we discuss the ungapped phases of the model.
3.1. Ungapped phases
Figure 2. Left: support of the spectral function in the ungapped phase A (small $\alpha$ large $\gamma$ phase) for a few values of $\alpha$. The dot at $z=-1/\gamma$ is the pole which is inside $\mathcal{C}$, while the other pole (not shown) is outside. Right: phase boundary between the ungapped phase A and the gapped phase.
Figure 3. Support of the spectral function (left panels) and the phase boundary between the gapped and ungapped phases (right panels). The panels from top to bottom correspond to ungapped phases B, C and D in the same order.
@@ 3.1. Ungapped phases
(30) $\sigma\left(-1+\dfrac{1}{1+\alpha z}+\dfrac{1}{1+\gamma z}\right)+\mathscr{N}=\mathscr{P}\oint_{\mathcal{C}}\frac{dz^{\prime}}{2\pi iz^{\prime}}\,\rho(z^{\prime})\,\frac{z+z^{\prime}}{z-z^{\prime}},\qquad z\in\mathcal{C},$
become closed contour integrals. Thus, we can immediately utilize elementary theorems to solve the model and compute various quantities of interest.
We assume the following ansatz based on the branch points of the potential ($z=-\frac{1}{\alpha},-\frac{1}{\gamma}$):
(31) $\rho(z)=\sum_{n=0}^{\infty}\frac{A_{n}}{z^{n}}+\sum_{n=0}^{\infty}\frac{B_{n}}{(z+1/\alpha)^{n}}+\sum_{n=0}^{\infty}\frac{C_{n}}{(z+1/\gamma)^{n}}.$
It is necessary to include a pole at $z=0$ in the ansatz to ensure that the normalization condition (26) is satisfied, and this pole must always lie inside $\mathcal{C}$. The coefficients in the ansatz are determined by substituting it into (30) and comparing both sides. In addition, the normalization condition may be used to determine any one coefficient.
There are four different solutions for $\rho(z)$ distinguished by whether the poles $z=-\frac{1}{\alpha}$ and $z=-\frac{1}{\gamma}$ are inside or outside $\mathcal{C}$. Each of these solutions corresponds to a distinct ungapped phase. Just as in the finite-$N$ case, the phases of the model are characterized by the branch points of the potential. There is, however, a critical difference between the finite-$N$ and large-$N$ transitions, which we will discuss at the end. We use the following nomenclature henceforth:
@@
(32) $\rho(z)=1+\sigma+\frac{\sigma}{1+\gamma z}-\frac{\sigma}{1+\alpha z}.$
The comparison also shows that the Lagrange multiplier $\mathscr{N}$ vanishes. With the spectral function known, the SU($N$) constraint (29) simplifies to
(33) $(1+\sigma)\ln{r_{0}}+\sigma\ln{\left(\frac{1-r_{0}\alpha}{r_{0}\gamma-1}\right)}+\sigma\ln\gamma=0$
@@
(34) $\alpha|z|^{1+1/\sigma}=\left|z+\frac{1}{\alpha}\right|^{-1}\,\left|z+\frac{1}{\gamma}\right|.$
As expected, $\mathcal{C}$ is not a unit circle in this phase. See the left panel in Fig. 2 for plots of $\mathcal{C}$ in phase A.
Figure 4. Schematic of the support $\mathcal{C}$ of the spectral function in the gapped phase (dashed, both panels). The dots on the negative real axis are the branch points of the potential ($z=-1/\alpha,-1/\gamma$). Solid closed curve in the left panel is the contour $\mathcal{C}^{*}$ used to solve integrals in the gapped phase. It can be deformed continuously to that shown in the right panel. See Fig. 5 for the actual contours in the gapped phase.
@@
(35) $\alpha=\frac{\gamma r_{0}-\sigma-1}{r_{0}(\gamma r_{0}(\sigma+1)-2\sigma-1)}$
@@
(36) $W_{n}:=\frac{1}{N}\langle\mathrm{tr}\,U^{n}\rangle=\oint_{C}\dfrac{dz}{2\pi iz}\rho(z)z^{n}=\begin{cases}(-1)^{n+1}\frac{\sigma}{\gamma^{n}}\hskip 42.67912ptn\geq 1\\ (-1)^{n+1}\sigma\alpha^{|n|}\hskip 34.14322ptn\leq-1\end{cases}$
If our model results from a lattice gauge theory, then these moments are related to the winding Wilson loops over a single plaquette [32]. We observe that
(37) $W_{n}\neq W_{-n}$
which indicates the presence of the sign problem. This behavior is completely expected, as our model’s potential is complex, in addition to having an SU($N$) geometry.
Lastly, we look at the free energy $F$, defined in (8). Since $F$ in the large-$N$ limit is just the rescaled effective action, $F=\frac{\sigma}{N}\mathcal{S}_{\text{eff}}$, the continuum version of (19) can be used to directly calculate the free energy. Another method is to first find the derivatives of the free energy with respect to each parameter of the model using
(38) $\frac{\partial F}{\partial t}=\left\langle\frac{\partial V(U)}{\partial t}\right\rangle=\oint_{C}\frac{dz}{2\pi iz}\rho(z)\frac{\partial V(z)}{\partial t},\qquad t=\sigma,\alpha,\gamma,$
@@
(39) $\frac{\partial F}{\partial\alpha}=\frac{\sigma^{2}}{\alpha-\gamma},\qquad\frac{\partial F}{\partial\gamma}=-\frac{\sigma^{2}}{\gamma}+\frac{\sigma^{2}}{\gamma-\alpha},\qquad\frac{\partial F}{\partial\sigma}=2\sigma\ln{\left(1-\frac{\alpha}{\gamma}\right)}.$
@@
(40) $F=\sigma^{2}\ln\left(1-\frac{\alpha}{\gamma}\right).$
In the tables below we summarize the important quantities in each phase. See Fig. 2 and Fig. 3 for plots of the support of the spectral function for each ungapped phase and the phase diagram near the ungapped-gapped transition.